OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 9 meters and base radius of 6 meters? Use three point one four for pi.

OpenStudy (wolf1728):

Cone Volume = (π • r² • h) / 3 Cone Volume = (3.14 * 3² * 9) / 3 Cone Volume = 84.78

OpenStudy (shiraz14):

@wolf1728 : You're correct